OpenStudy (anonymous):

Solve 4x^2 = 12x + 40. Show Work

OpenStudy (anonymous):

4x^2 = 12x + 40 x^2 = 3x + 10 x^2-3x-10 = 0 on solving u will get x = 5 or -2